The equation -12x2 + 4800 = 0 represents the path of a softball being pitched. The solutions represent the amount of time the softball is in the air. Solve the equation using square roots and determine which solution is correct in this real world solution.


1
Expert's answer
2021-02-23T12:42:54-0500

-12x² + 4800 = 0

=> 12x² = 4800

x² = "\frac{4800}{12} = 400"

x = "\sqrt{400} = ±20"


x is either +20 or -20


Since x represents the amount of time spent by the softball in air, and we're dealing with real world solution, and in real world solution, time can never be negative. That implies that x is +20


x = 20 is the correct solution in this real world solution